Aptitude Test No. of claims settled in a month


65 68

70 60

60 62

77 80

75 85

50 40

53 52

73 62

65 60

82 81


Using regression analysis on the above data answer the following question:

A) if a new employee Sarah scores 88 on the test how many claims is she like to settle?

B) Interpret the coefficient of determination


1
Expert's answer
2021-02-24T07:10:26-0500

Answer

The mean is given by

"\bar{x}"= "1 \over n""\u03a3x"

Applied test

"\bar{x}" = 67


Claims

"\bar{y}"= 65

The slope is given by


β1= "Σ (x-\bar{x}) (y-\bar{y}) \over Σ (x-\bar{x})^2"


β1= "1149 \over 976"= 1.17725


The intercept is given by


 β0= "\bar{y}"β1"\bar{x}"


β0= 65-67x 1.17725 = -13.876


The regression line is given by

y=-13.876+ 1.17725


a)


Excel formula




The output is given below




From the output, the regression equation is 

No. of claims settled in a month =-13.879 + 1.177254 (Aptitude Test)


To find the number of settlements substitute Aptitude Test = 88 in the regression equation.

The regression equation is 

No. of claims settled in a month =-13.879 + 1.177254 (Aptitude Test)

substitute Aptitude Test = 88

No. of claims settled in a month =-13.879 + 1.177254 (88)

=89.72

No. of claims settled in a month =90 (approximately)

Thus, the number of claims is she like to settle is 90.


b)

Error sum of square is given by


ESS = "Σ (y-\bar{y})^2" = 399.335


Total sum of square is given by


TSS= "Σ (y-\bar{y})^2" = 1752


The coefficient of determination is given by


R2=1-"ESS \over TSS"



R2=1-"399.335 \over 1752"


It shows that around 77.21% variation in the number of claims settled is explained by the aptitude test scores.